kaylak:

I have 4 math questions IDK

kaylak:

@Vocaloid

kaylak:

1 attachment
kaylak:

@Vocaloid

kaylak:

@Shadow

iGreen:

Are you familiar with this? \(\sf\dfrac{sin(A)}{a}=\dfrac{sin(B)}{b}=\dfrac{sin(C)}{c}\)

kaylak:

kinda yes

iGreen:

Okay, we can apply it here to find 'x'.

kaylak:

is the answer b?

iGreen:



iGreen:

Can you set up the equation?

iGreen:

We can use: \(\sf\dfrac{sin(A)}{a}=\dfrac{sin(B)}{b}\) Plug in the given information.

iGreen:

You there?

kaylak:

yes sorry

iGreen:

Okay, can you plug them in?

iGreen:

A = 42 a = x B = 180 - (61 + 42) b = 13

kaylak:

sorry my parents keep calling me

kaylak:

B is 77

kaylak:

i think its b

kaylak:

@iGreen

kaylak:

hello

kaylak:

@Vocaloid

iGreen:

Yeah, angle B is 77, so plug all the values into this: \(\sf\dfrac{sin(A)}{a}=\dfrac{sin(B)}{b}\)